OpenStudy (dottiecat):

What is the solution to the equation 3/4(4c+16)=2c+9? c= ___

OpenStudy (dottiecat):

@serenity2001

OpenStudy (thenaughtyserperior):

Is that 3/4 as a fraction or 3 divided by 4?

OpenStudy (dottiecat):

3/4 as a fraction

OpenStudy (thenaughtyserperior):

Okay, so distribute 3/4 into (4c + 16). So multiply both 4c and 16.

OpenStudy (dottiecat):

20c

OpenStudy (dottiecat):

nvm c = -3 @TheNaughtySerperior

OpenStudy (serenity2001):

3/4 (4c + 16) = 2c + 9 3(4c + 16) / 4 = 2c + 9 3(c + 4) = 2c + 9 3c + 12 = 2c + 9 -12 -12 3c = 2c - 3 -2c -2c c = -3

OpenStudy (dottiecat):

Yeah... I figured it out already....

OpenStudy (dottiecat):

\[\frac{-2n-2 }{ 3 } = 12? what is n